OpenStudy (anonymous):

4x−3(5x−8)=23−9(x+2)

OpenStudy (anonymous):

4x-15x+40=23-9x-18....-11x+40=5-9x

OpenStudy (anonymous):

...-11x+9x=5-40...-2x=-35...x=35/2

OpenStudy (anonymous):

...sorry one moment, the answer is x=19/2

OpenStudy (anonymous):

its okay am just checking the steps on mine to see where i messed up at

OpenStudy (anonymous):

the +40 should be a +24

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

tks Mathty

OpenStudy (anonymous):

your welcome :), if you have any questions let me know

OpenStudy (anonymous):

i have a few more but i hate to bother you

OpenStudy (anonymous):

by all means ask away, I'll see what I can do

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x-5(4-9x)\ge-3(6x+11)\]

OpenStudy (anonymous):

x-20+45x (greater or equal) -18x-33... 46x+18x (greater or equal) -33+20... x\[\ge\]-13/64

OpenStudy (anonymous):

gtg though, good luck

OpenStudy (anonymous):

tks